OpenStudy (anonymous):

What is the quotient (91y^3 + 21y^2 – 35y) ÷ 7y

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

13y^2 + 3y – 5

OpenStudy (lgbasallote):

\[\frac{91y^3 + 21y^2 - 35y}{7y} \implies \frac{91y^3}{7y} + \frac{21y^2}{7y} - \frac{35y}{7y}\]